the manufacturer of a bag of sweets clams that there is a 90% chance that the bag contains some toffees. if 20 bags are chosen, what is the probability that 

a) All the bags contain toffees   

b) More than 18 bags contain toffees 

 


1
Expert's answer
2021-11-11T18:55:27-0500

Let "X=" the number of the bags which contain some toffees: "X\sim Bin(n, p)."

Given "n=20, p=0.9, q=1-p=1-0.9=0.1."

a)


"P(X=20)=\dbinom{20}{20}(0.9)^{20}(0.1)^{20-20}"

"\approx0.1215766546"

b)


"P(X>18)=P(X=19)+P(X=20)"

"=\dbinom{20}{19}(0.9)^{19}(0.1)^{20-19}+\dbinom{20}{20}(0.9)^{20}(0.1)^{20-20}"

"\approx0.2701703435+0.1215766546"

"\approx0.391747"